### v1

# A generalized Apollonius problem

Abstract.
We present a self-contained solution of a generalized Apollonius problem of constructing a circle that intersects each of the three pair-wise non-intersecting circles at one of the three given angles.
1. Introduction
In the $3^{\text{rd}}$ century BC, Apollonius of Perga posted and solved the famous problem of constructing – with the help of a straightadge and a compass only – a circle tangent to three given circles - see, for example, [4, 2, 3] for a general statement.
The significance of this problem can hardly be overstated - no wonder the title of the Presidential address delivered in 1967 by the great geometer of the $20^{\text{th}}$ century H.S.M. Coxeter to the Canadian Mathematical Congress at the joint meeting with the Mathematical Association of America was ”The Problem of Apollonius.” At the beginning of his presentation Coxeter stated, ”As Sir Thomas Heath remarks [6, p. 182], this problem ’has exercised the ingenuity of many distinguished geometers, including Vieta and Newton.’” (The results presented in this address were later published in [2].)
As a result, there are many ingenious approaches to the Apollonius problem. Unfortunately, Apollonius’ solution of his problem has not survived. However, François Viète found a solution by considering limiting cases, where any of the three given circles can be reduced to a point or blown up into a line, which one may believe to be the Apollonius’ approach.
The method of inversion yields an efficient way of solving Apollonius problem, where two of the three given circles are transformed either into two parallel lines or into two concentric circles. This approach, in part, utilizes the fact that inversion preserves angles between two intersecting circles. We use this fact, together with the availability of an inversion that transforms two non-intersecting circles into concentric ones, to solve the problem of constructing – using a straightedge and a compass only – a cirle that intersects each of the three given circles at one of the three given angles from the interval $[0,90^{0}]$. The Apollonius problem is a particular case of this problem where the three given angles are zeros.
To solve this problem, we first assume that such a circle exists and perform an inversion of the four circles that transforms two of the three given circles into concentric ones. Then we find a formula for a possible radius of the image of the circle in question and formulas for possible distances from its center to the centers of the images of the given circles. Finally, using these three formulas, we construct a circle $\mathcal{C}^{\prime}$ that does intersect each of the images of the three given ones at one of the three given angles. As the square of an inversion is the identity map, the inversion of $\mathcal{C}^{\prime}$ delivers a circle that intersects each of the three given circles at one of the three given angles.
2. Preliminaries
2.1. The (acute) angle between two circles
Let $A$ be a point of intersection of two circles as in Figure 1. Then the acute angle at which these circles intersect is defined as the acute angle $\alpha$ between the tangent lines $t_{1}$ and $t_{2}$ to the circles at the point $A$. Clearly this angle coincides with the acute angle between the lines that pass through $O_{1}$ and $A$ and $O_{2}$ and $A$:
Figure 1
2.2. Constructibility
Let us define what it means for an interval, or an angle, to be constructible.
Definition 2.1.
Fix a unit interval, and let $a_{1},a_{2},\dots,a_{n}$ be intervals. Denote their lengths by the same letters. An interval of length
(1) $a=f(a_{1},a_{2},\dots,a_{n})$
is called constructible if it can be drawn with the help of straightedge and a compass. In this case, the number $a$ is also called constructible. An angle $\alpha$ is called constructible if $\cos\alpha$ (equivalently, $\sin\alpha$) is a constructible number. If an interval (angle) can be constructed avoiding the use of the unit interval, we will call it properly constructible or p-consructible.
We presume that the reader is familiar with the following straightedge and compass p-constructions:
  1. (a)
    Finding the mid-point of a given interval.
  2. (b)
    Drawing the line through a given point and perpendicular to a given line.
  3. (c)
    Drawing the line through a given point and parallel to a given line.
Example 2.1.
P-construct the circle that passes through three non-collinear points $A$, $B$, and $C$.
Solution: Utilizing constructions (a) and (b), draw perpendicular bisectors of the intervals $[AB]$ and $[BC]$. The point of intersection of these bisectors is the center of the circle though $A$, $B$, and $C$.
To outline the scope of constructible intervals it is sufficient to consider a few examples where we use the constructions (a) - (c) without explicit references.
Example 2.2.
Given intervals $a$ and $b$, p-construct an interval of length $x=\sqrt{ab}$.
Solution: Draw a segment $[AB]$ and a point $C\in[AB]$ such that $|AC|=a$ and $|CB|=b$ as in Figure 2. Then draw a semi-circle with $[AB]$ as its diameter. Next, construct the line through $C$ perpendicular to the line $AB$. If $X$ is the point of intersection of this line with the semi-circle, then $\angle AXB=90^{0}$, implying that the triangles $ACX$ and $XCB$ are similar. Therefore, denoting $|CX|=x$, we have
$\frac{|CX|}{|AC|}=\frac{|CB|}{|CX|}\ \ \Longleftrightarrow\ \ \frac{x}{a}=\frac{b}{x}\ \ \Longleftrightarrow\ \ x=\sqrt{ab},$
thus, the interval $[CX]$ has the desired length.
Figure 2
Remark 2.1.
In particular, if $b=1$, then $x=\sqrt{a}$, that is, square roots of constructible intervals are also constructible.
Example 2.3.
Given intervals $a$, $b$, and $c$, p-construct an interval of length $x=\displaystyle\frac{ab}{c}$.
Solution: Referring to Figure 3, draw segments $[OB]$ and $[OC]$ and the point $A\in[OC]$ such that $|OA|=a$, $|OC|=c$, and $|OB|=b$. Next, construct the line through $A$ parallel to $BC$, and let $X$ be the point of intersection of this line with $OB$. Then, denoting $|OX|=x$, we have
$\frac{|OX|}{|OA|}=\frac{|OB|}{|OC|}\ \ \Longleftrightarrow\ \ \frac{x}{a}=\frac{b}{c}\ \ \Longleftrightarrow\ \ x=\frac{ab}{c}.$
Therefore, the interval $[CX]$ has the desired length.
Figure 3
Remark 2.2.
In particular, when $c=1$ ($b=1$), we have $x=ab$ (respectively, $x=\displaystyle\frac{a}{c}$ ), that is, the products or ratios of constructible intervals are constructible.
Example 2.4.
Given intervals $a$, $b$, $c$, $d$, and an angle $\alpha\in(0,\pi)$, construct an interval of length
$e=\frac{\sqrt{a^{2}+bc\sin\alpha}}{\sqrt[4]{d}}.$
Solution: In view of Remark 2.2, one can construct an interval of length $a^{2}$ and an interval of length $bc$, hence an interval of length $bc\sin\alpha$. Thus, an interval of length $x=a^{2}+bc\sin\alpha$ can be constructed. Next, in view of Remark 2.1, one can construct an interval of length $\sqrt{x}=\sqrt{a^{2}+bc\sin\alpha}$. Further, as $4]{d}=\sqrt{\sqrt{d}}$, Remark 2.1 yields an interval of length $y=\sqrt[4]{d}$. Finally applying Remark 2.2, we end up with an interval of length $\frac{\sqrt{x}}{y}=e$.
2.3. Criterion of Constructibility
We can now state that if the function $f$ in (1) can be written as involving only the four arithmetic operations and operations of finding square roots, then the interval $a$ is constructible. In fact, it can be proven, as it is done in [1, pp. 129-134], that the converse statement is also true.
3. Statement of the Problem
3.1. Apollonius Three Circle Problem
Referring to Figure 4, we can state the famous Apollonius problem as follows:
Problem 3.1.
Given three circles $\mathcal{C}_{1}$, $\mathcal{C}_{2}$, $\mathcal{C}_{3}$, p-construct a circle $\mathcal{C}$ that is tangent to each of them, that is, intersects each at one point.
Figure 4
Note that, generically, there are eight different solutions to Apollonius problem, as the circle in question may touch one, or two, or all three given circles from outside. Moreover, the circles $\mathcal{C}_{1}$, $\mathcal{C}_{2}$, $\mathcal{C}_{3}$ may not be positioned as conveniently as in Figure 4.
Referring to Figure 5, we propose the following, that is a generalization of the Apollonius problem:
Problem 3.2.
Given three circles  $\mathcal{C}_{1}$, $\mathcal{C}_{2}$, $\mathcal{C}_{3}$ with centers at $O_{1}$, $O_{2}$, $O_{3}$ and radii $r_{1}$, $r_{2}$, $r_{3}$, respectively, p-construct a circle $\mathcal{C}$ that intersects $\mathcal{C}_{1}$ at an angle $0^{0}\leq\alpha\leq 90^{0}$, $\mathcal{C}_{2}$ at an angle $0^{0}\leq\beta\leq 90^{0}$, and $\mathcal{C}_{3}$ at an angle $0^{0}\leq\gamma\leq 90^{0}$.
To avoid technicalities, we assume that the circles  $\mathcal{C}_{1},\mathcal{C}_{2},\mathcal{C}_{3}$ are proper, their centers $O_{1}$, $O_{2}$, $O_{3}$ form a proper triangle, and that
$|O_{1}O_{2}|\geq r_{1}+r_{2},\ \ |O_{1}O_{3}|\geq r_{1}+r_{3},\text{ \ and\ }|O_{2}O_{3}|\geq r_{2}+r_{3},$
as in Figure 5:
Figure 5
4. Solution of Problem 3.2
An elegant solution of the Apollonius problem - in a particular case given by Figure 4 - is given via an application of inversion as follows. First, the radii of the three given circles are enlarged by the same interval - so that two of the resulting circles become tangent at, say, a point $O$ and then performing an inversion with the center $O$ resulting in these two circles turn into two parallel lines.
Our solution is based on the following facts:
  1. (A)
    An inversion preserves the angles between two circles - see [3, Theorem 5.51].
  2. (B)
    Given two non-intersecting circles $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$, there exists a circle $\mathcal{C}_{\mathcal{I}}$ such that $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ turn into concentric circles under the inversion with respect to $\mathcal{C}_{\mathcal{I}}$ see [3, Theorem 5.71]. Note that this property can be used to find more solutions to the Apollonius problem - see [5].
4.1. Inversion and its Properties
Referring to Figure 6, consider a circle with center $O$ and radius $R$, which we will call the inversion circle and denote $\mathcal{C}_{\mathcal{I}}$.
Figure 6
Points $P\neq O$ and $P^{\prime}\neq O$ are called the inversion images of each other with respect to $\mathcal{C}_{\mathcal{I}}$ if $P$ and $P^{\prime}$ lie on the same ray aminating from $O$ and
$|OP|\cdot|OP^{\prime}|=R^{2}.$
This transformation of the plane - minus the point $O$ is called the inversion with respect to $\mathcal{C}_{\mathcal{I}}$. If we denote this inversion by $\mathcal{I}$, then $\mathcal{I}(P)=P^{\prime}$, $\mathcal{I}(P^{\prime})=P$, hence $\mathcal{I}^{2}(P)=P$ for every $P\neq O$, and $\mathcal{I}(P)=P$ whenever $P\in\mathcal{C}_{\mathcal{I}}$.
Figure 6 gives a way of p-constructing the point $P^{\prime}$ if $P$ is given and vise versa: If $P$ is given, let $A$ be the intersection of $\mathcal{C}_{\mathcal{I}}$ with the circle for which $[OP]$ is a diameter. Then $PA$ is tangent to $\mathcal{C}_{\mathcal{I}}$, and if $AP^{\prime}\perp OP$, the triangles $AOP$ and $P^{\prime}OA$ are similar, so we have
$\frac{|OP|}{R}=\frac{R}{|OP^{\prime}|}\ \Longleftrightarrow\ |OP|\cdot|OP^{\prime}|=R^{2}.$
If $P^{\prime}$ is given, let $A\in\mathcal{C}_{\mathcal{I}}$ be such that $AP^{\prime}\perp OP^{\prime}$. Then, if $P\in OP^{\prime}$ is such that $AP\perp OA$, it follows as above that $|OP|\cdot|OP^{\prime}|=R^{2}$.
Next, it can be proven - see [3, Theorem 5.45] - that
  1. (i)
    If a circle $\mathcal{C}$ does not pass through the center $O$ of $\mathcal{C}_{\mathcal{I}}$, then $\mathcal{I}(\mathcal{C})$ is a circle.
  2. (ii)
    If a circle $\mathcal{C}$ passes through $O$, then $\mathcal{I}(\mathcal{C})$ is a straigh line.
Let us now describe a construction that will need to be performed repeatedly in the sequel without explicit reference:
Example 4.1.
Given a circle $\mathcal{C}$ that does not pass through the center of an inversion circle $\mathcal{C}_{\mathcal{I}}$, p-construct $\mathcal{I}(\mathcal{C})$.
Solution: Pick three points $A,B,C$ on $\mathcal{C}$ and p-construct the points $\mathcal{I}(A)$, $\mathcal{I}(B)$, $\mathcal{I}(C)$ using the procedure given by Figure 6. Then p-construct the circle $\mathcal{I}(\mathcal{C})$ as passing through the points $\mathcal{I}(A)$, $\mathcal{I}(B)$, $\mathcal{I}(C)$ as in Example 2.1.
4.2. P-constructing a circle $\mathcal{C}_{\mathcal{I}}$ in (B)
Referring to [3, Ch. 5, § 7] and Figure 7, we can describe the process of p-constructing possible centers for the circle $\mathcal{C}_{\mathcal{I}}$ from the statement (B) above as follows.
Figure 7
Let $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ be non-intersecting circles with centers at $O_{1}$ and $O_{2}$ and $radii$ $r_{1}$ and $r_{2}$, respectively. P-constructing a right triangle with the hypotenuse $[O_{1}O_{2}]$ and one of the sides of length $|r_{1}-r_{2}|$, we can draw the common outer tangents to the circles $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$. Denote by $A_{1}$, $B_{1}$ and $A_{2}$, $B_{2}$ the points of intersection of these tangents with circles $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$, respectively, and let $M_{1}$ and $M_{2}$ be the mid-points of segments $A_{1}B_{1}$ and $A_{2}B_{2}$, respectively. Next, denote by $C$ the point of intersection of the lines $O_{1}O_{2}$ and $M_{1}M_{2}$.
P-construct a point $D\in\mathcal{C}_{1}$, as we did in Figure 6 with point $A$, such that $CD$ is tangent to $\mathcal{C}_{1}$ and draw the circle of radius $|CD|$ with the center at $C$. Let this circle intersect the segment $[O_{1}O_{2}]$ at points $\mathcal{O}_{\mathcal{I}_{1}}$ and $\mathcal{O}_{\mathcal{I}_{2}}$ counting from $O_{1}$. According to [3], each of the points $\mathcal{O}_{\mathcal{I}_{1}}$ and $\mathcal{O}_{\mathcal{I}_{2}}$ can serve as the center of the inversion cirle in (B).
For the reader’s convenience, Figure 8 complements Figure 5 by depicting the circle $\mathcal{C}_{\mathcal{I}}$, in green, that we will use to ensure that $\mathcal{C}_{2}^{\prime}=\mathcal{I}(\mathcal{C}_{2})$ and $\mathcal{C}_{3}^{\prime}=\mathcal{I}(\mathcal{C}_{3})$ are concentric. Note that the result of this inversion, given in Figure 9, lies entirely inside $\mathcal{C}_{\mathcal{I}}$.
Figure 8
Assume that there exists a circle $\mathcal{C}$ that satisfies the conditions in Problem 3.2, as in Figure 5. Let us apply the inversion $\mathcal{I}$ given by the circle $\mathcal{C}_{\mathcal{I}}$ in Figure 8 to the four circles in Figure 5. In view of the statement (A) above, the result looks as depicted in Figure 9.
Figure 9
Referring to Figure 9, we apply the Law of Cosines to triangles $\mathcal{O}^{\prime}A\mathcal{O}_{2}^{\prime}$ and $\mathcal{O}^{\prime}B\mathcal{O}_{2}^{\prime}$:
$(r^{\prime})^{2}+(r_{2}^{\prime})^{2}+2r^{\prime}r_{2}^{\prime}\cos\beta=(r^{\prime})^{2}+(r_{3}^{\prime})^{2}-2r^{\prime}r_{3}^{\prime}\cos\gamma,$
implying that
(2) $2r^{\prime}(r_{2}^{\prime}\cos\beta+r_{3}^{\prime}\cos\gamma)=(r_{3}^{\prime})^{2}-(r_{2}^{\prime})^{2}.$
We also calculate
(3) $|\mathcal{O}^{\prime}\mathcal{O}_{1}^{\prime}|=\sqrt{(r^{\prime})^{2}+(r_{1}^{\prime})^{2}+2r^{\prime}r_{1}^{\prime}\cos\alpha}$
and
(4) $|\mathcal{O}^{\prime}\mathcal{O}_{2}^{\prime}|=\sqrt{(r^{\prime})^{2}+(r_{2}^{\prime})^{2}+2r^{\prime}r_{2}^{\prime}\cos\beta}.$
Now we are ready to solve Problem 3.2:
4.3. P-constructing a circle $\mathcal{C}$
Assume first that at least one of the angles $\alpha,\beta,\gamma$ is less than $90^{0}$. Without loss of generality, let it be $\beta$ and/or $\gamma$. Then we can consider a circle $\mathcal{C}^{\prime}$ with radius
(5) $r^{\prime}=\frac{(r_{3}^{\prime})^{2}-(r_{2}^{\prime})^{2}}{2(r_{2}^{\prime}\cos\beta+r_{3}^{\prime}\cos\gamma)}=\frac{r_{3}^{\prime}r_{3}^{\prime}}{2(r_{2}^{\prime}\cos\beta+r_{3}^{\prime}\cos\gamma)}-\frac{r_{2}^{\prime}r_{2}^{\prime}}{2(r_{2}^{\prime}\cos\beta+r_{3}^{\prime}\cos\gamma)}$
and the center $\mathcal{O}^{\prime}$ such that $|\mathcal{O}^{\prime}\mathcal{O}_{1}^{\prime}|$ and $|\mathcal{O}^{\prime}\mathcal{O}_{2}^{\prime}|$ are given by (3) and (4), respectively, as depicted in Figure 10.
Note that, in view of Problem 2.3, the decomposition in (5) implies, as the interval $c=2(r_{2}^{\prime}\cos\beta+r_{3}^{\prime}\cos\gamma)$ is clearly p-constructible, that $r^{\prime}$ is p-constructible. Then, p-constructing a triangle with sides $r^{\prime}$ and $r_{1}^{\prime}$ and the angle $180^{0}-\alpha$ between them, we see that $|\mathcal{O}^{\prime}\mathcal{O}_{1}^{\prime}|$, and similarly $|\mathcal{O}^{\prime}\mathcal{O}_{2}^{\prime}|$, are p-constructible intervals. Hence, we can p-consruct a circle $\mathcal{C}^{\prime}$.
Figure 10
Now, referring to Figure 10 and applying the Law of Cosines to triangle $\mathcal{O}^{\prime}A\mathcal{O}_{1}^{\prime}$, we obtain
$(r^{\prime})^{2}+(r_{1}^{\prime})^{2}-2r^{\prime}r_{1}^{\prime}\cos\alpha^{\prime}=(r^{\prime})^{2}+(r_{1}^{\prime})^{2}+2r^{\prime}r_{1}^{\prime}\cos\alpha,$
implying that
(6) $\cos\alpha^{\prime}=-\cos\alpha,\text{ \ hence \ }\alpha^{\prime}=180^{0}-\alpha.$
Similarly, considering triangle $\mathcal{O}B\mathcal{O}_{2}^{\prime}$, we infer
(7) $\cos\beta^{\prime}=-\cos\beta,\text{ \ hence \ }\beta^{\prime}=180^{0}-\beta.$
Finally, the Law of Cosines for triangle $\mathcal{O}^{\prime}C\mathcal{O}_{2}^{\prime}$ together with (2) yield
$2r^{\prime}(r_{2}^{\prime}\cos\beta+r_{3}^{\prime}\cos\gamma^{\prime})=(r_{3}^{\prime})^{2}-(r_{2}^{\prime})^{2}=2r^{\prime}(r_{2}^{\prime}\cos\beta+r_{3}^{\prime}\cos\gamma),$
so
(8) $\cos\gamma^{\prime}=\cos\gamma,\text{ \ hence \ }\gamma^{\prime}=\gamma.$
[…]
Let now
$\alpha=\beta=\gamma=90^{0}.$
If we apply (p-construct) the inversion $\mathcal{I}$, then $\mathcal{C}^{\prime}=\mathcal{I}(\mathcal{C})$ must intersect each of the concentric circles $\mathcal{C}_{2}^{\prime}=\mathcal{I}(\mathcal{C}_{2})$ and $\mathcal{C}_{3}^{\prime}=\mathcal{I}(\mathcal{C}_{3})$ at a $90$-degree angle. This implies that $\mathcal{I}(\mathcal{C})$ is a line, since there is no circle that intersects each of the two concentric circles at a $90$-degree angle - in such a case, equation (2) would yield $r_{2}^{\prime}=r_{3}^{\prime}$, which is never the case.
Note that, in view of the properties (i) and (ii) in Subsection 4.1, this implies that if a circle $\mathcal{C}$ intersects each of the two non-intersecting circles $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$, as in Figure 7, at a $90$-degree angle, then it passes through the center of $\mathcal{C}_{\mathcal{I}}$ (actually, it passes through both $O_{1}$ and $O_{2}$ in Figure 7). This is a striking fact by itself showcasing the remarkable usefulness of the inversion.
Observe that, being orthogonal to each of the circles  $\mathcal{C}_{1}^{\prime}=\mathcal{I}(\mathcal{C}_{1})$, $\mathcal{C}_{2}^{\prime}=\mathcal{I}(\mathcal{C}_{2})$ (and $\mathcal{C}_{3}^{\prime}=\mathcal{I}(\mathcal{C}_{3})$), the line $\mathcal{I}(\mathcal{C})$ must pass through the centers $\mathcal{O}_{1}^{\prime}$ and $\mathcal{O}_{2}^{\prime}=\mathcal{O}_{3}^{\prime}$, so it can be drawn immediately. Costructing the inversion $\mathcal{I}$ of this line, we arrive at a circle $\mathcal{C}$ that intersects each of the circles  $\mathcal{C}_{1}$, $\mathcal{C}_{2}$, $\mathcal{C}_{3}$ at a $90$-degree angle.
Remark 4.1.
As with the classical Apollonius problem, there is more than one solution to Problem 3.2. For example, one can see that changing """""-"" - " for "+""""+"" + " in front of $2r^{\prime}r_{1}^{\prime}\cos\alpha$ in (3) will not alter the values of $\alpha,\beta,\gamma$. Therefore, such a change is admissible. The reader can find more admissible changes in definitions of $r^{\prime}$, $\mathcal{O}^{\prime}\mathcal{O}_{1}^{\prime}$, and $\mathcal{O}^{\prime}\mathcal{O}_{2}^{\prime}$ resulting in different solutions of Problem 3.2.